# Bounds on the photon mass via the Shapiro effect in the solar system

### v1

@@ Appendix A The massless case: Eddington’s $\gamma$ parameter and the PPN metric
In the main text we determined the gravitational time delay in the case of a massive photon that follows the geodesic path from emitter to receiver and back. This approach is often replaced by a simplified calculation using a straight line to approximate the true geodesic. In this appendix we would like to highlight the difference between the two, in particular for the standard case of a massless photon.
Our focus in this paper is to obtain limits on the photon mass by comparing the predictions with a finite mass to observational data. For this, it is useful to parametrize deviations from general relativity and the parametrized post-Newtonian (PPN) metric does just that Poisson ; Will_book ; Will_exp . The most relevant modification is that characterized by Eddington’s $\gamma$ parameter, which appears in the metric coefficients, cf. eq. (39), as ($\mathcal{M}=GM_{\odot}/c^{2}$)
$A(r)\approx 1-\frac{2\mathcal{M}}{r}\quad{\rm and}\quad B(r)\approx 1+\gamma\frac{2\mathcal{M}}{r}\,.$ (29)
In general relativity $\gamma=1$ and deviations may signal effects of new physics.
@@ A.1 Straight-line approximation
In the literature, most of the discussions of the Shapiro effect is done in terms of the usual Schhwarzschild coordinates using a straight line as an approximation for the curved path light takes. Let us constrain the motion to the $x-y$ plane, where emitter and receiver are at $(-x_{e},r_{0})$ and $(x_{r},r_{0})$, respectively, with $r_{0}$ being the point of closest approach.
The photon moves in the $x$-direction and the line element is $ds^{2}=0=A(r)c^{2}dt^{2}-B(r)dx^{2}$, so that, to leading order in $\mathcal{M}/r$,
@@ A.2 Geodesic solution
with $A(r),B(r)$ given by eq. (29).
Restricting the motion to the equatorial plane $\theta=\pi/2$, we may define a new Lagrangean
@@
We may now contrast the PPN expressions for a massless photon obtained in the straight-line approximation, cf. eq. (31), with that determined via the geodesic, cf. eq. (37). Clearly, the Newtonian travel time is the same in both, as it should; the post-Newtonian terms, however, are different. In particular, for $\gamma=1$ we have a difference of $8GM_{\odot}/c^{3}\approx 40\,{\rm\mu s}$ between eqs. (31) and (37) that is purely due to the paths followed. Being practically constant (independent of $r_{0}$), this difference is not measurable in observations with varying perihelion distances, as was the case of the Cassini experiment Cassini_1 ; Cassini_2 , cf. Sec. III.3. It does, however, contribute to the total coordinate time elapsed, so we will employ eq. (37) as the standard PPN expression for the Shapiro effect in Sec. III.4.
@@ Appendix B Geodesic equations in isotropic coordinates
Let us start by looking at eq. (40c) and consider movement in the equatorial plane $\theta=\pi/2$ with $\dot{\theta}=0$. These choices clearly satisfy this equation and imply that the trajectory will remain on this plane for all time, so we shall limit ourselves to this case. With this choice, the equations above reduce to
@@
If we move away from the Sun we have $A(r)\rightarrow 1$, so that eq. (42a) becomes $2c^{2}\dot{t}=\epsilon$, but now we are in flat Minkowski spacetime, where $\dot{t}=\gamma$, the Lorentz boost factor (not to be confused with Eddington’s $\gamma$ parameter), thus $2\gamma c^{2}=\epsilon$. We are therefore led to state that $\epsilon=2E/m_{\gamma}$ with $E$ being the total energy of the massive photon. Moving on to eq. (42d), at large distances, we get $B(r)\rightarrow 1$, so $-2r^{2}\dot{\phi}=\Lambda$. We recognize that, if $\Lambda=-2\ell/m_{\gamma}$, we recover the Newtonian expression for the (orbital) angular momentum, $\ell$.
@@
Note that the equations above are independent of the constants of motion $\epsilon$ and $\Lambda$.
@@ Appendix C Deflection angle in the straight-line approximation
In Sec. II.1 we based our calculation on the actual geodesics followed by a massive, ultra-relativistic particle when passing near a massive body. In this appendix we repeat the calculations under the assumption that the geodesics may be approximated by straight lines. For simplicity, we take the trajectory to lie on the $x-y$ plane. In this case we have $M_{\odot}$ at $y=0$ and the trajectory is taken to be parallel to the $y$-axis with $dx^{\beta}\approx(c/v,0,1,0)dy$, where $v$ is the velocity of the massive particle with 4-momentum $p^{\alpha}=(E/c,0,p^{y},0)$ Nieves .
Here we follow the approach from Refs. Ohanian ; Nieves and assume that the massive photon is moving along the positive $y$-direction with impact parameter $b$ along the $x$-axis. To first order in $GM_{\odot}/bc^{2}$, the total change in the massive photon’s 4-momentum is
@@
Energy and momentum are conserved and we are evaluating the deflection in asymptotically flat regions, where the metric is Minkowskian and the dispersion relation $E^{2}=m_{\gamma}^{2}c^{4}+(cp^{y})^{2}$ holds. Furthermore, energy and velocity are related via $E=m_{\gamma}c^{2}/\sqrt{1-v^{2}/c^{2}}$. Since $m_{\gamma}c^{2}\ll E$, we have

### v2

@@ Appendix A The massless case: Eddington’s $\gamma$ parameter and the PPN metric
In the main text we determined the gravitational time delay in the case of a massive photon following the geodesic path from emitter to receiver and back. This approach is often replaced by a simplified calculation using a straight line to approximate the true geodesic. Here we highlight the difference between the two, in particular for the standard case of a massless photon.
Given the many different alternative theories of gravitation, it is generally useful to devise a convenient parametrization to quantify possible deviations from general relativity – the parametrized post-Newtonian (PPN) metric does just that Poisson ; Will_book ; Will_exp . The most relevant modification is that characterized by Eddington’s $\gamma$ parameter, which appears in the metric coefficients as
$g_{00}\approx 1-\frac{2\mathcal{M}}{r}\quad{\rm and}\quad g_{ii}\approx 1+\gamma\frac{2\mathcal{M}}{r}\,,$ (31)
where $\mathcal{M}=GM_{\odot}/c^{2}$ and the index $i$ runs over the spatial coordinates; in general relativity $\gamma=1$.
@@ A.1 Straight-line approximation
In the literature, most of the discussions of the Shapiro effect is done in terms of the usual Schhwarzschild coordinates using a straight line as an approximation for the curved path light takes. Let us constrain the motion to the $x-y$ plane, where emitter and receiver are at $(-x_{e},r_{0})$ and $(x_{r},r_{0})$, respectively, with $r_{0}$ being the point of closest approach of the light ray.
The photon moves in the $x$-direction and the line element is $ds^{2}=0=g_{00}c^{2}dt^{2}-g_{xx}dx^{2}$, so that, to leading order in $\mathcal{M}/r$,
@@ A.2 Geodesic solution
with $A(r)=g_{00}$ and $B(r)=g_{ii}$ given by eq. (31). Restricting the motion to the equatorial plane $\theta=\pi/2$, we may define a new Lagrangean
@@
We may now contrast the PPN expressions for a massless photon obtained in the straight-line approximation, cf. eq. (33), with that determined via the geodesic, cf. eq. (39). Clearly, the Newtonian travel time is the same in both, as it should; the post-Newtonian terms, however, are different. In particular, for $\gamma=1$ we have a difference of $8GM_{\odot}/c^{3}\approx 40\,{\rm\mu s}$ between eqs. (33) and (39) that is purely due to the paths followed.
@@ Appendix B Geodesic equations in isotropic coordinates
Let us start with eq. (42c) and consider movement in the equatorial plane $\theta=\pi/2$ with $\dot{\theta}=0$. These choices satisfy this equation and imply that the trajectory will remain on this plane for all time. The equations above then reduce to
@@
If we move away from the Sun we have $A(r)\rightarrow 1$, so that eq. (44a) becomes $2c^{2}\dot{t}=\epsilon$, but now we are in flat Minkowski spacetime, where $\dot{t}=\gamma$, the Lorentz boost factor (not to be confused with Eddington’s $\gamma$ parameter), thus $2\gamma c^{2}=\epsilon$. We may then say that $\epsilon=2E/m_{\gamma}$ with $E$ being the total energy of the massive photon. Moving on to eq. (44d), at large distances $B(r)\rightarrow 1$, so $-2r^{2}\dot{\phi}=\Lambda$; if $\Lambda=-2\ell/m_{\gamma}$ we recover the Newtonian expression for the (orbital) angular momentum, $\ell$.
@@ Appendix C Deflection angle in the straight-line approximation
In Sec. II.1 we based our calculation on the geodesics followed by a massive, ultra-relativistic particle passing near a massive body. Here we repeat the calculations under the assumption that the geodesics may be approximated by straight lines. For simplicity, we take the trajectory to lie on the $x-y$ plane. In this case we have $M_{\odot}$ at $y=0$ and the trajectory is taken to be parallel to the $y$-axis with $dx^{\beta}\approx(c/v,0,1,0)dy$, where $v$ is the velocity of the massive particle with 4-momentum $p^{\alpha}=(E/c,0,p^{y},0)$ Nieves .
Following Refs. Ohanian ; Nieves we assume that the massive photon is moving along the positive $y$-direction with impact parameter $b$ along the $x$-axis. To first order in $GM_{\odot}/bc^{2}$, the total change in its 4-momentum is
@@
Energy and momentum are conserved and we are evaluating the deflection in asymptotically flat regions, where the dispersion relation $E^{2}=m_{\gamma}^{2}c^{4}+(cp^{y})^{2}$ holds. Furthermore, energy and velocity are related via $E=m_{\gamma}c^{2}/\sqrt{1-v^{2}/c^{2}}$. Since $m_{\gamma}c^{2}\ll E$, we have